# SSAP: A Shape-Sensitive Adversarial Patch for Comprehensive Disruption of Monocular Depth Estimation in Autonomous Navigation Applications

### v1

@@ I Introduction
Previous efforts in patch-based adversarial attacks, which focused solely on CNN-based MDE [1, 2], produced relatively weak adversarial patches. These patches had a limited impact on the depth estimation of specific objects such as vehicles and pedestrians, with their influence confined to the overlapping area between the patch and the input image. There is significant potential for improvement in expanding the sphere of influence of these patches.
@@ II-E Penalized Depth Loss
$L_{d_{1}}=(\left|d_{t}-d_{adv}\right|\odot M_{P})$ (7)
$L_{d_{2}}=\left|d_{t}-d_{adv}\right|\odot(M_{f}-M_{P})$ (8)
@@ V Related Work
Unlike existing physical attacks that have targeted tasks like object detection [27, 28], image classification [29, 30], and face recognition [31, 32], the area of attacks on depth estimation has received relatively limited attention. Zhang [33] proposes an attack designed to enhance performance in a universal attack scenario. Wong [34] introduces a strategy for generating targeted adversarial perturbations, which then randomly alter the associated depth map. It’s worth noting that these two attacks center around digital perturbations, rendering them less suitable for real-world applications.

### v2

@@ I Introduction
Previous efforts in patch-based adversarial attacks, which focused solely on CNN-based MDE [1, 2, 18], produced relatively weak adversarial patches. These patches had a limited impact on the depth estimation of specific objects such as vehicles and pedestrians, with their influence confined to the overlapping area between the patch and the input image. There is significant potential for improvement in expanding the sphere of influence of these patches.
@@ II-E Penalized Depth Loss
$L_{d_{1}}=\frac{1}{m\times n}\sum_{i,j}\left|d_{t}-d_{adv}\right|\odot M_{P}$ (7)
$L_{d_{2}}=\frac{1}{m\times n}\sum_{i,j}\left|d_{t}-d_{adv}\right|\odot(M_{f}-M_{P})$ (8)
@@ V Related Work
Unlike existing physical attacks that have targeted tasks like object detection [28, 29, 30], image classification [31, 32, 33], and face recognition [34, 35], the area of attacks on depth estimation has received relatively limited attention.
@@ VI Conclusion
Acknowledgment
This research was partially funded by the NYUAD Center for Cyber Security (CCS), funded by Tamkeen under the NYUAD Research Institute Award G1104 and the Technology Innovation Institute (TII) under the ”CASTLE: Cross-Layer Security for Machine Learning Systems IoT” project.